106. Eleanor scores 680 on the mathematics part of the SAT. The distribution of SAT math scores in recent years has been Normal with mean 504 and standard deviation Gerald takes the ACT Assessment mathematics test and scores 27. ACT math scores are Normally distributed with mean 23.8 and standard deviation 3.8 a. What is Elanors standardized score? b. What is Geralds standardized score? c. Assuming that both tests measure the same kind of ability, who has the higher score? Round to 3 decimal places. Round to 3 decimal places. Gerald. Elanor. They both did equally well.please answer all parts
Answer #1

as we know that standardized score =(X-mean)/std deviation

a)

Elanor's standardized score =(680-504)/106=1.660

Gerald's standardized score =(27-23.8)/3.8=0.842

Elanor has higher score having higher z value